OpenStudy (anonymous):

** CORRECT ME PLEASE ! [WORK IS SHOWN] Addison earned a gross income of $42,750 last year. She made $537.56 in IRA contributions, donated $918 to her favorite charity and paid $1,351.35 in home mortgage interest. If Addison claims a standard deduction of $5,700 and her exemption is $3,650, what is her taxable income? 5700 + 3650 + 537.56 + 918 + 1351.35 = 12,156.91. 42,750.00 - 12,156.91 = 30,593.09 ?<---taxable income

OpenStudy (anonymous):

$30,593.09 $36,293.09 $31,511.09 $32,862.44

OpenStudy (radar):

I am assuming that you are permitted to use standard deduction and also itemize the deductions listed ( I am not a tax lawyer), then your work appears proper and the taxable income is as you say, and it agrees with the first choice.

OpenStudy (anonymous):

@radar :c my teacher told me it was wrong ?

OpenStudy (radar):

Well the math looked ok, maybe some of those items used to diminish the gross income are nort permitted.

OpenStudy (anonymous):

I got the answer !

OpenStudy (ivettef365):

what is the answer

OpenStudy (anonymous):

32862.44

OpenStudy (radar):

Also the final value is normally rounded off so that there is no cents. Maybe it is $30,593, but all the answers shows some pennies. I am perplexed.......if you put the 5700 back on the answer36,293.09 which is the suggestion of ivettef365

OpenStudy (anonymous):

What I did ! 5700 + 3650 =9350 + 537.56 = 9887.56 - [32862.44 ] you don't add the deductions i forgot that part !

OpenStudy (anonymous):

I mean 9887.56 - 42750.0 = 32862.44

OpenStudy (anonymous):

In questions like these all you have to do is take the gross income - IRA contributions - standard deduction - exemption so you're right

OpenStudy (radar):

As I said I am not a tax lawyer, and right now the IRS is in disfavor lol

OpenStudy (radar):

Thanks @murrcat so it was a matter on knowing what could be subtracted "legally" from the gross income.

OpenStudy (radar):

Sorry NaCl (salt) but I am afraid I wasn't much help.